OpenStudy (anonymous):

6x+9/3 x 2x^2/16x-12 simplify

OpenStudy (blurbendy):

\[\frac{ 6x + 9 }{ 3 } * \frac{ 2x^2 }{ 16x - 12 }\]

OpenStudy (blurbendy):

does it look like that?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

you can simplify it to this first: \[\frac{ 2x^2(6x+9) }{ 3(16x-12 }\] the three on the bottom comes from dividing by 3 in the left equation. Then you can distribute the 2x^2 on the top, and the 3 on the bottom